OpenStudy (anonymous):

Suppose f(x) is a function with continuous first and second derivatives on the closed interval [0,4] whose values for f and f' at x=0 and x=4 are given as x=0; f(x)=3; f'(x)=3 x=4; f(x)=5; f'(x)=-1 Prove there exists a value of d, 0

OpenStudy (anonymous):

isnt this suppose 2 b in da algebra section

OpenStudy (anonymous):

no. I believe it has to do with the Mean Value Theorem or Rolle's Theorem.

OpenStudy (anonymous):

it is mvt directly

OpenStudy (anonymous):

\[f'(0)=-1,f'(4)=3\] \[\frac{f'(4)-f'(0)}{4-0}=\frac{-1-3}{4}=-1\]